# Almost device-independent certification of GME states with minimal measurements

### 28 Feb 2024 v1

@@ III.3 $N$-qubit generalized $W$ states
for every $j$, where ${B^{\prime\prime}}$ denotes Bob’s auxiliary system and the state $|\psi_{N}\rangle$ is given by,
where $|\xi\rangle_{\mathbf{B}^{\prime\prime}E}\in\mathcal{H}_{B_{1}^{\prime\prime}}\otimes\ldots\otimes\mathcal{H}_{B_{N-1}^{\prime\prime}}\otimes\mathcal{H}_{E}$ denotes the auxiliary state.
The reader will find the proof for the above theorem in Appendix C.
IV Discussion
In our work, we demonstrated that although it’s very difficult to create self-testing methods with the fewest measurements possible, having trust in one of the parties makes it possible to create 1SDI schemes for many types of GME states. We focused on graph states, Schmidt states, and generalized $W$ states. Another important finding is that we can certify a pair of operators that are mutually unbiased with minimal entanglement across any number of parties in one process.
@@ IV Discussion
Our work raises several follow-up questions. An obvious one is how to expand our approach to certify more types of genuinely entangled multipartite states. In fact, it would be interesting to see whether the assumption considered here that one of the measurement device is trusted allows one to design certification scheme for any GME state based on the minimal number of measurements per observes. A related question that arises here is whether the ideas presented above can be combined with the results of Ref [44, 45] to design self-testing schemes for multipartite states which are based on finite statistics. What is more, it would have been interesting to explore whether the 1SDI certification scheme for graph states of any local dimension provided here could be generalized to stabilizer subspaces (cf. Ref. [46]). At the same time, as far as the experimental testing of our schemes is concerned, it is crucial to understand how resilient they are to experimental errors. Additionally, we can use our self-testing results for secure multipartite randomness generation.
V Acknowledgments
@@
This project was funded within the QuantERA II Programme (VERIqTAS project) that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733 and from the Polish National Science Center (project No 2021/03/Y/ST2/00175). Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101080086 NeQST.
Supplementary Material: Almost device-independent certification of GME states with minimal measurements
Appendix A Appendix A: Self-testing any Graph state
@@
[…]
@@
$\left(\mathbbm{1}_{AE}\otimes\bigotimes_{i=1}^{N-1}U_{j}\right)|\psi_{N}\rangle_{A\mathbf{B}E}=|G\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (49)
@@
[…]
@@
Notice, that terms inside the large brackets in both the above formulas commute. Now, by using the fact that $\omega^{\gamma_{0,k}}Z_{A}^{-\gamma_{0,k}}X_{A}^{-1}=X_{A}^{-1}Z_{A}^{-m_{e,k0}}$, we can multiply (56) by $\omega^{\gamma_{0,k}}$ and subtract (57) to obtain that
@@
$U_{j}\,B_{j,0}\,U_{j}^{\dagger}=Z_{B^{\prime}_{j}}\otimes\mathbbm{1}_{B^{\prime\prime}_{j}},\qquad U_{j}\,B_{j,1}\,U_{j}^{\dagger}=X_{B^{\prime}_{j}}\otimes\mathbbm{1}_{B^{\prime\prime}_{j}}\qquad\forall j.$ (63)
@@
$\forall i\qquad\overline{S}(G)_{i}|\psi_{N}\rangle=\left(\bigotimes_{i=1}^{N-1}U_{j}^{\dagger}\right)S(G)_{d,i,AB_{1}^{\prime\prime}\ldots B_{N}^{\prime\prime}}\otimes\mathbbm{1}_{B_{1}^{\prime\prime}\ldots B_{N}^{\prime\prime}E}\left(\bigotimes_{i=1}^{N-1}U_{j}\right)|\psi_{N}\rangle=|\psi_{N}\rangle,$ (64)
@@
where $S(G)_{d,i}$ are the stabilizing operators of a graph state $|G\rangle$ of the form (11). As discussed above, the graph state $|G\rangle$ is the unique eigenstate of such operators with eigenvalue $1$. Consequently, we have that
$\left(\bigotimes_{i=1}^{N-1}U_{j}\right)|\psi_{N}\rangle_{A\mathbf{B}E}=|G\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (65)
@@
where $|\xi\rangle_{\mathbf{B}^{\prime\prime}E}\in\bigotimes_{j=1}^{N-1}\mathcal{H}_{B_{j}^{\prime\prime}}\otimes\mathcal{H}_{E}$ denotes the auxiliary state. ∎
Appendix B Appendix B: Self-testing Schmidt states

### 13 Mar 2026 v2

@@ III.3 $N$-qubit generalized $W$ states
for every $i$, where $B^{\prime\prime}$} denotes Bob’s auxiliary system and the state $|\psi\rangle_{A\mathbf{B}E}$ is given by,
$\left(\mathbbm{1}_{AE}\otimes\bigotimes_{i=1}^{N-1}U_{i}\right)|\psi\rangle_{A\mathbf{B}E}=|\psi_{W}(\boldsymbol{\alpha})\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (33)
where $|\xi\rangle_{\mathbf{B}^{\prime\prime}E}$ denotes the auxiliary state.
As before, the proof of this statement is deferred to Appendix C.
Let us notice that self-testing methods for the standard $N$-qubit $W$ state (i.e., one for which $\alpha_{l}=1/\sqrt{N}$ for every $l=1,\ldots,N$) were introduced in Refs. [16, 17]. Here, by making a mild assumption that a single party is trusted, we can significantly generalise these results to a multi- parameter class of genuinely entangled states.
IV Extending to device-independence
We now show how our 1SDI certification methods for multipartite quantum states can be extended to the device-independent regime, restricting for simplicity to the case where the quantum states are locally qubits. To this end, we combine our quantum steering scenario with the network-assisted approach presented in Ref. [20]. More precisely, we introduce an extra party Charlie in between Alice and Bobs and also an additional source of quantum particles distributing a state $|\phi\rangle_{AC}$ to Alice and Charlie (cf. Fig. 1). The aim of Charlie is to play the CHSH game with Alice in order to certify her measurements and also to teleport the local state $\rho_{A}=\mathrm{Tr}_{\mathbf{B}}\rho_{A\mathbf{B}}$ of the multipartite state $\rho_{A\mathbf{B}}$ to Alice. Charlie performs three measurements (cf. Fig. 3), each having two outcomes. The correlations observed by all parties are described by $\{p(a,c,b_{1},\ldots,b_{N-1}|x,z,y_{1},\ldots,y_{N-1})\}$ where $z\in\{0,1,2\}$ and $c\in\{0,1\}$ denote the input and output of Charlie, respectively. The first two Charlie’s measurements labeled by $z=0,1$ are performed on Charlie’s subsystem of $|\phi_{AC}\rangle$, whereas the third one on both Charlie’s subsystems. Let us now go through the steps in details.
Figure 3: Extended Bell scenario. An additional party, Charlie, enforces trust in Alice by certifying his measurements. In the considered scenario there are two sources of states: one in between Alice and Charlie distributing bipartite states $|\phi\rangle_{AC}$ and one in between Charlie and N1N-1 Bob distributing NN partite states.
Step I: Certifying Alice’s observables
To certify Alice’s observables, Alice and Charlie simply play the Clauser-Horne-Shimony-Holt (CHSH) game [51] on the state distrubted by the bipartite source. If the correlations observed by them $\{p(a,c|x,y)\}$ violate the CHSH Bell inequality maximally, that is, they achieve the maximal quantum value of the functional
$I_{\mathrm{CHSH}}:=\langle A_{0}C_{0}\rangle+\langle A_{0}C_{1}\rangle+\langle A_{1}C_{0}\rangle-\langle A_{1}C_{1}\rangle,$ (34)
which is $I_{\mathrm{CHSH}}=2\sqrt{2}$, then they infer that there exist local unitary operations $U_{A}$ and $U_{C}$ such that
$U_{A}\otimes U_{C}|\psi\rangle_{AC}=|\phi^{+}\rangle_{A^{\prime}C^{\prime}}\otimes|\xi\rangle_{A^{\prime\prime}C^{\prime\prime}},$ (35)
where $|\xi\rangle_{A^{\prime\prime}C^{\prime\prime}}$ is some auxiliary state, and
$U_{A}\,A_{0}\,U_{A}^{\dagger}=Z\otimes\mathbbm{1}_{A^{\prime\prime}},\quad U_{B}\,A_{1}\,U_{B}^{\dagger}=X\otimes\mathbbm{1}_{B^{\prime\prime}}.$ (36)
Thus, Alice, by playing the CHSH game with Charlie, can certify her measurements to be, up to the above equivalences, the two Pauli observables which are used in the steering inequality (31).
Step II: Teleporting $\rho_{A}$ Charlie to Alice
Once the state between Alice and Charlie is certified to be maximally entangled the local state $\rho_{A}$ of the multipartite state $\rho_{A\mathbf{B}}$ from Charlie to Alice using the third Charlie’s measurement corresponding to $z=2$, which is performed on both systems received by Charlie. We only restrict our analysis to the first outcome of the Bell measurement, as no communication is allowed between Alice and Bob during the protocol. The reason for this teleportation step rather than directly sending the state $\rho_{A}$ to Alice is that the measurements of Alice are only certified on the local support of $\mathrm{Tr}_{C}|\phi\rangle\!\langle\phi|_{AC}$. If the local support is different, then the measurement cannot be guaranteed [see [31]]. Consequently, in the next step, we are only interested in the correlations $\{p(a,0,b_{1},\ldots,b_{N}|x,2,y_{1},\ldots,y_{N})\}$, keeping in mind the self-tested state and measurements in the previous step.
Step-III: Self-testing with certified measurements
In the last step Alice together with all Bobs can now use the 1SDI certification methods developed in the previous section to certify a given multipartite state in $\rho_{A\mathbf{B}}$. However, now the trusted observables $Z$ and $X$ must be replaced by those given in (36), which have been previously certified by Alice and Charlie.
Consider now an extended Bell scenario operator $\overline{\mathcal{S}}$ constructed, for instance, from the steering operator (III.3) in which $Z$ and $X$ have been replaced by the measurements $A_{1},A_{2}$ certified (35). It is not difficult to observe that the new steering operator can be written as
$\overline{\mathcal{S}}=U_{A}^{\dagger}(\mathcal{S}\otimes\mathbbm{1}_{A^{\prime\prime}})U_{A},$ (37)
where $\mathcal{S}$ are any of the above-introduced steering operators. Consequently, the LHS and quantum bounds remain the same in the extended Bell and the standard steering scenario. Moreover, if $\langle\overline{\mathcal{S}}\rangle=\beta_{Q}$, then we readily have that $\langle\mathcal{S}\otimes\mathbbm{1}_{A^{\prime\prime}}\rangle=\beta_{Q}$. Consequently, any state $|\psi_{A\mathbf{B}}\rangle$ attaining the maximal value of $\overline{\mathcal{S}}$ must also satisfy $\mathrm{Tr}(\mathcal{S}\rho_{A^{\prime}\mathbf{B}})=\beta_{Q}$ where $\rho_{A^{\prime}\mathbf{B}}=\mathrm{Tr}_{A^{\prime\prime}}(U_{A}\psi_{A\mathbf{B}})$. This is the exact condition that we utilise in the above 1SDI certification proofs, and thus it can be straightforwardly used here to obtain that
$U_{A}\otimes U_{\mathbf{B}}|\psi_{A^{\prime}\mathbf{B}E}\rangle=|\tilde{\psi}_{A^{\prime}\mathbf{B^{\prime}}}\rangle\otimes|\xi_{\mathbf{B^{\prime\prime}}E}\rangle$ (38)
where $|\psi^{\prime}\rangle_{A\mathbf{B}}$ is the ideal state and the system $E$ is added to purify $\rho_{A^{\prime}\mathbf{B}}$. Consequently, the state generated by the source is certified to be $U_{A}\otimes U_{\mathbf{B}}|\psi_{A\mathbf{B}E}\rangle=|\tilde{\psi}_{A^{\prime}\mathbf{B^{\prime}}}\rangle\otimes|\xi_{A^{\prime\prime}\mathbf{B^{\prime\prime}}E}\rangle$.
V Discussion
In our work, we showed that while designing self-testing methods with a minimal number of measurements is highly challenging, introducing trust in one of the parties enables the construction of 1SDI schemes for a broad class of genuinely multipartite entangled (GME) states. We focused on graph states, Schmidt states, and generalised $W$ states. Another important finding is that using an arbitrarily low amount of entanglement it is possible to certify, in a single process, a pair of mutually unbiased operators for an arbitrary number of observers. In inspired by the results of Ref. [20], we also lifted the 1SDI certification to DI self-testing by including an additional party to certify the measurement of the trusted party. While in this work we have only considered the case when the trusted party performs two-outcome measurements, this approach technique can be straightforwardly generalised to a higher number of outcomes with the trusted party, provided a suitable Bell inequality exists which allows one to self-test the measurements of the trusted party. One can also utilise the network-based self-testing of any measurement presented in [21] to certify arbitrary Alice’s measurements; however, such an approach will be highly complicated and difficult to implement. We believe that our results open a path towards DI self-testing strategies based on the minimal number of measurements per party which are applicable to any genuinely entangled state. At the moment no such a scheme exists.
@@ V Discussion
Our work raises several follow-up questions. An obvious one is how to expand our approach to certify other classes of genuinely entangled multipartite states. In fact, it would be interesting to see whether the assumption considered here that one or few of the measurement devices are trusted allows one to design certification scheme for any GME state based on the minimal number of measurements per observes. A related question that arises here is whether the ideas presented above can be combined with the results of Ref. [52, 53] to design self-testing schemes for multipartite states which are based on finite statistics. What is more, it would have been interesting to explore whether the 1SDI certification scheme for graph states of any local dimension provided here could be generalized to stabilizer subspaces (cf. Ref. [54]). At the same time, as far as the experimental testing of our schemes is concerned, it is crucial to understand how resilient they are to experimental errors. Additionally, we can use our self-testing results for secure multipartite randomness generation.
VI Acknowledgments
@@
This project was funded within the QuantERA II Programme (VERIqTAS project) that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733 and from the Polish National Science Center (project No 2021/03/Y/ST2/00175). Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. SS also acknowledges the National Science Centre, Poland, grant Opus 25, 2023/49/B/ST2/02468. RA also acknowledges received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101080086 NeQST.
Appendix A Self-testing any graph state
@@
Assume that the steering inequality (41) is maximally violated when the $0^{\text{th}}$ party Alice chooses the observables $A_{i}$ defined to be $A_{0}=Z$, $A_{1}=X$ and all the other parties choose their observables as $B_{i,j}$ for $i=1,2\ldots,N-1$ and $j\in\{0,1\}$ acting on $\mathcal{H}_{B_{i}}$. Let us say that the state which attains this violation is given by $|\psi_{N}\rangle_{A\mathbf{B}}\in\mathbbm{C}^{d}\otimes\bigotimes_{i=1}^{N-1}\mathcal{H}_{B_{i}}\otimes\mathcal{H}_{E}$, then the following statement holds true for any $d$: $\mathcal{H}_{B_{i}}=\mathcal{H}_{B^{\prime}_{i}}\otimes\mathcal{H}_{B^{\prime\prime}_{i}}$, where $\mathcal{H}_{B^{\prime}_{i}}\equiv\mathbbm{C}^{d}$, $\mathcal{H}_{B^{\prime\prime}_{i}}$ is some finite-dimensional Hilbert space, and there exists a local unitary transformation on Bob’s side $U_{i}:\mathcal{H}_{B_{i}}\rightarrow\mathcal{H}_{B_{i}}$, such that
$\forall j,\quad U_{i}\,B_{j,0}\,U_{i}^{\dagger}=Z_{B^{\prime}_{i}}\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},\qquad U_{i}\,B_{j,1}\,U_{i}^{\dagger}=X_{B^{\prime}_{i}}\otimes\mathbbm{1}_{B^{\prime\prime}_{i}}$ (53)
@@
$\left(\mathbbm{1}_{AE}\otimes\bigotimes_{i=1}^{N-1}U_{i}\right)|\psi_{N}\rangle_{A\mathbf{B}E}=|G\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (54)
where $|\xi\rangle_{\mathbf{B}^{\prime\prime}E}\in\bigotimes_{i=1}^{N-1}\mathcal{H}_{B_{i}^{\prime\prime}}\otimes\mathcal{H}_{E}$ denotes the auxiliary state.
@@
Let us begin by considering the first term and second term of the steering operator (40) corresponding to the stabilizers when choosing the $0^{\text{th}}$ vertex and vertices connected to it respectively. As concluded in Eq. (52), we have that
@@
Notice, that terms inside the large brackets in both the above formulas commute. Now, by using the fact that $\omega^{\gamma_{0,k}}Z_{A}^{-\gamma_{0,k}}X_{A}^{-1}=X_{A}^{-1}Z_{A}^{-\gamma_{0,k}}$, we can multiply (61) by $\omega^{\gamma_{0,k}}$ and subtract (62) to obtain that
@@
$U_{i}\,B_{i,0}\,U_{i}^{\dagger}=Z_{B^{\prime}_{i}}\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},\qquad U_{i}\,B_{i,1}\,U_{i}^{\dagger}=X_{B^{\prime}_{i}}\otimes\mathbbm{1}_{B^{\prime\prime}_{i}}\qquad\forall i.$ (68)
@@
$\forall i\qquad\overline{S}(G)_{i}|\psi_{N}\rangle=\left(\bigotimes_{i=1}^{N-1}U_{i}^{\dagger}\right)S(G)_{d,i,AB_{1}^{\prime\prime}\ldots B_{N}^{\prime\prime}}\otimes\mathbbm{1}_{B_{1}^{\prime\prime}\ldots B_{N}^{\prime\prime}E}\left(\bigotimes_{i=1}^{N-1}U_{i}\right)|\psi_{N}\rangle=|\psi_{N}\rangle,$ (69)
@@
where $S(G)_{d,i}$ are the stabilizing operators of a graph state $|G\rangle$ of the form (13). As discussed above, the graph state $|G\rangle$ is the unique eigenstate of such operators with eigenvalue $1$. Consequently, we have
$\left(\bigotimes_{i=1}^{N-1}U_{i}\right)|\psi_{N}\rangle_{A\mathbf{B}E}=|G\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (70)
@@
where $|\xi\rangle_{\mathbf{B}^{\prime\prime}E}\in\bigotimes_{i=1}^{N-1}\mathcal{H}_{B_{i}^{\prime\prime}}\otimes\mathcal{H}_{E}$ denotes the auxiliary state. ∎
Let us now restrict ourselves to $d=2$ and find the robust self-testing statement. Before proceeding, let us state a general Lemma which will be required for the robust proof.
Lemma 1.
[…]
where $|\psi^{\prime}\rangle=U|\psi\rangle$, and for simplicity, we have omitted all the identities acting on the second subsystem $B$.
The proof of this lemma can be found in Appendix B of Ref. [55]. Let us now derive the robustness statement.
Theorem 2.
Let us consider a graph state $|G\rangle$ that we want to certify robustly. For this graph, we consider a hierarchy of nodes as: (0) $n(0)$: all nodes connected to the $0$-th node (trusted node), (1) $n_{1}(j)$: all nodes connected to some $j$-th such that $j\in n(0)$, which are not neighbors of the node 0, nor the node 0 itself, (2) $n_{2}(j)$: all nodes connected to some node in $n_{1}(j)$ such that they are not connected to 0 and $n(j)$, etc. Since we consider graphs which are not disjoint, for each vertex there exists a path connecting it to the trusted node.
Assume now that the steering inequality (41) attains a value close to the maximal violation $\langle\mathcal{I}_{2}(G,N)\rangle\geq\beta_{Q}-\varepsilon$, when the $0^{\text{th}}$ party (trusted) Alice chooses the observables $A_{i}$ defined to be $A_{0}=Z$, $A_{1}=X$. Then, the state $|\psi\rangle_{A\mathbf{B}}$ and the observables $B_{i,j}$ for $i=1,\ldots,N-1$ and $j=0,1$ are certified as
[…]
where $g=\sum_{k}n_{k}a_{k}$ is the number of nodes connected in the $k$th heirarchy and $a_{k}=(8/3)(4^{k+1}-1)$ and for node $i$ belonging to $k$th heirarchy,
[…]
where $|\psi^{\prime}\rangle=\bigotimes_{i=1}^{N-1}U_{{B_{i}}}|\psi\rangle$ and we have dropped the indices $A\mathbf{B}E$.
Proof.
Let us consider the steering operator $\mathcal{I}_{2}(G,N)$ given in Eq. (40) which in the particular case of $d=2$ simplifies to
$\mathcal{I}_{2}(G,N)$ $=$ $2\left[X_{A}\otimes\bigotimes_{j\in n(0)}B_{j,0}+\sum_{j\in n(0)}Z_{A}\otimes B_{j,1}\otimes\bigotimes_{j^{\prime}\in n(j)\setminus\{0\}}B_{j^{\prime},0}+\sum_{j\notin n(0)\cup{0}}B_{j,1}\otimes\bigotimes_{j^{\prime}\in n(j)}B_{j^{\prime},0}\right],$ (74)
where the factor $2$ is a consequence of the fact that all local observables are hermitian in this case.
Let us then observe that it admits the following sum-of-squares decomposition, $\beta_{Q}\mathbbm{1}-\mathcal{I}_{2}(G,N)=\sum_{j=1}^{N}P_{j}P_{j}^{\dagger}$ where the operators $P_{j}$ are given by
$P_{0}$ $=$ $\mathbbm{1}-X_{A}\otimes\bigotimes_{i\in n(0)}B_{i,0},$
$P_{j}$ $=$ $\mathbbm{1}-Z_{A}\otimes B_{j,1}\otimes\bigotimes_{j^{\prime}\in n(j)\setminus\{0\}}B_{j^{\prime},0}\quad\mathrm{for}\quad j\in n(0),$
$P_{j}$ $=$ $\mathbbm{1}-B_{j,1}\otimes\bigotimes_{j^{\prime}\in n(j)}B_{j^{\prime},0}\quad\mathrm{for}\quad j\notin n(0)\cup{0}.$ (75)
[…]
Let us now choose the kthk-th vertex such that $k\in n(0)$, take two of the above relations and expand using (A) as
[…]
Now, using the fact that $X,Z$ are hermitian and unitary for $d=2$ and that the vector norm is unitarily invariant, the above inequalities imply that
[…]
Exploiting the unitary invariance of the vector norm once more we can multiply the expressions under the norm by $Z_{A}$ and $X_{A}$, respectively, to obtain
[…]
Let us now consider the second inequality in (77) and by using the unitary invariance of the norm as well as the fact that all observables are unitary, let us expand it to the following form
[…]
Then, we the aid of the triangle inequality for the vector norm, one shows that the first inequality in (A) and (79) imply that
[…]
Analogously, from by suitably modifying the first inequality in (77) and combining it with the the second inequality in (A) through the triangle inequality, one can obtain
[…]
In deriving these inequalities we also exploited the fact that $B_{j,0}$ commute among each other if they act on different sites.
Now, using the fact that $Z_{A}X_{A}=-X_{A}Z_{A}$, and then exploiting the triangle inequality, we have that
[…]
To all the observables $B_{k,i}$ for $k\in n(0)$ and $i=0,1$ (which are those that are measured by the neighbors of the trusted party) we can apply Lemma 1 (where for simplicity we additionally assume that $B_{k,i}$ decompose under Jordan’s lemma only into nontrivial $2\times 2$ blocks), to conclude that there exists a local unitary operation $U_{i}$ acting on $i$th node such that
[…]
where $|\psi^{\prime}\rangle=\bigotimes_{i\in n(0)}U_{{i}}|\psi\rangle$.
The above steps allow us to find the robust bounds to observables for the nodes $j$ directly connected to the trusted party $j\in n(0)$. Based on the relations in Eq. (83), we can follow the same reasoning as above to prove analogous relations for the observables measured by the neighbors of all nodes in $n(0)$ that do not themselves belong to $n(0)$ and are not the trusted party. To this aim, let us now consider one of such nodes $j\in n(0)$ and prove that relations similar to (83) hold true for all nodes $k\notin n(0)\cup\{0\}$ connected to $j$. We can follow the same steps as above using the SOS $P_{j}$ and $P_{k}$ (A) to get
[…]
As the observables $B_{j,0},B_{j,1}$ for $j\in n(0)$ are certified as in Eq. (83), this allows us to obtain
[…]
Following the same steps from Eqs. (A) to (82), we obtain
[…]
This allows us to conclude from Lemma 1 that for each $i\in n(j),i\notin n(0)\cup\{0\}$, there exists a unitary operation $U_{i}$ such that
[…]
where $|\psi^{\prime\prime}\rangle=\bigotimes_{i\in n(j)},i\notin n(0)\cup\{0\}U_{{i}}|\psi\rangle$.
We now define an hierarchy of nodes as: (0) $n(0)$: all nodes connected to the $0$-th node (trusted node), (1) $n_{1}(j)$: all nodes connected to some $j$-th such that $j\in n(0)$, which are not neighbors of the node 0, nor the node 0 itself, (2) $n_{2}(j)$: all nodes connected to some node in $n_{1}(j)$ such that they are not connected to 0 and $n(j)$, etc. We proceed in this way until all vertices are taken care of; clearly, since we consider graphs which are not disjoint, for each vertex there exists a path connecting it to the trusted node. At each level the error in the relations (83) and (86) for the observables $B_{i,1}$ adds up as $8,40,168,\ldots$; one can check that this series is described $a_{k}=(8/3)(4^{k+1}-1)$ for $k$-th level of the hierarchy. More precisely, Bob’s observables satisfy
[…]
Let us again consider the steering operator $\mathcal{I}_{2}(G,N)$ (40) and observe that $2\beta_{Q}[\beta_{Q}\mathbbm{1}-\mathcal{I}_{2}(G,N)]=(\beta_{Q}\mathbbm{1}-\mathcal{I}_{2}(G,N))^{2}$ where we used the fact that $\mathcal{I}_{2}(G,N)^{2}=\beta_{Q}\mathbbm{1}$ where $\beta_{Q}=2N$. Consequently, from the assumption that $\langle\mathcal{I}_{2}(G,N)\rangle\geq\beta_{Q}-\varepsilon$, we infer the following inequality,
[…]
As $\mathcal{I}_{2}(G,N)$ is composed of the observables $B_{i,0},B_{i,1}$, which are certified as (83), we can rewrite the above formula using the triangle inequality as
[…]
where $\mathcal{I}_{2}^{\mathrm{id}}=2\sum_{i=1}^{N}\overline{S}(G)_{2,i}$ [c.f. Eq. (13)] and $\mu=2\sqrt{N}+g$ such that the factor $g$ takes into account the graph structure such that the errors add up based on the hierarchy of nodes defined below Eq. (87). Suppose that at $k$th heirarchy, there are $n_{k}$ nodes, then $g=\sum_{k}n_{k}a_{k}$.
Consider now the eigensystem of the ideal steering operator, $\mathcal{I}_{2}^{\mathrm{id}}|G_{l}\rangle=\lambda_{l}|G_{l}\rangle$. It is simple to observe that the difference between the highest and second-highest eigenvalues is $4$ as the second-highest eigenvalue is attained by $|G_{l}\rangle$ where $l_{k}=1$ for some $k$ and $l_{i}=0$ for all $i$ such that $i\neq k$. Let us now express the above formula (90) using the eigensystem of $\mathcal{I}_{l}^{\mathrm{id}}$ as
[…]
Let us now implement the following expansion of the state $|\psi^{\prime}\rangle$
$|\psi^{\prime}\rangle=\sum_{i=1}^{2^{N}}\alpha_{i}|G_{i}\rangle|\xi_{i}\rangle,$ (92)
where $|G_{i}\rangle$ are the eigenstates of the ideal steering operator $\mathcal{I}^{\mathrm{id}}$,i.e., $\mathcal{I}^{\mathrm{id}}|G_{i}\rangle=\lambda_{i}|G_{i}\rangle$. Then, $\alpha_{i}$ are nonnegative coefficient such that $\sum_{i}\alpha_{i}^{2}=1$, and $|\xi_{i}\rangle$ are some auxiliary vectors $|\xi_{i}\rangle\in\bigotimes_{i=1}^{N}\mathcal{H}_{B^{\prime\prime}_{i}}$. The fact that $\alpha_{i}$ can be taken nonnegative stems from the freedom in choosing the auxiliary vectors $|\xi_{i}\rangle$. Let us also enumerate the eigenvalues $\lambda_{i}$ so that $\lambda_{1}=2N$ which is the quantum bound $\beta_{Q}$; then, the corresponding eigenstate $|G_{1}\rangle$ is the graph state to be self-tested, i.e., $|G_{1}\rangle\equiv|G\rangle$.
Using this decomposition, the expression (91) allows one to deduce that
[…]
which on expanding the left-hand side results in
$\sum_{l=1}^{2^{N}}\alpha_{l}^{2}(\beta_{Q}-\lambda_{l})^{2}\leq\mu^{2}\varepsilon.$ (94)
Let us now consider the eigenvalues $\lambda_{i}$ of the ideal steering operator $\mathcal{I}^{\mathrm{id}}$. Clearly, the largest eigenvalue is exactly the quantum bound $\beta_{Q}=2N$ and therefore the corresponding term in the above sum vanishes. It is then not difficult to observe that the next largest eigenvalue is 2N42N-4, which allows us to write the following inequality
$\sum_{l=1}^{2^{N}}\alpha_{l}^{2}(\beta_{Q}-\lambda_{l})^{2}\geq 4\sum_{l=2}^{2N}\alpha_{l}^{2},$ (95)
where we used the fact that $(\beta_{Q}-\lambda_{l})^{2}\geq 4$ for every $l=2,\ldots,2^{N}$. Consequently, we have the following inequality
$\sum_{\begin{subarray}{c}l=2\end{subarray}}^{2^{N}}\alpha_{l}^{2}\leq\mu^{2}\frac{\varepsilon}{4}.$ (96)
We then use the fact that $\sum_{i}\alpha_{i}^{2}=1$ which together with (96) implies that $\alpha_{1}\geq\alpha_{1}^{2}\geq 1-(18N)^{2}\frac{\varepsilon}{4}$. Thus,
$\langle{\psi^{\prime}}|(|G\rangle|\xi_{1}\rangle)=\alpha_{1}\geq 1-\mu^{2}\frac{\varepsilon}{4},$ (97)
which directly implies that
[…]
which is the desired inequality (72). This completes the proof.
Appendix B Self-testing Schmidt states